OpenStudy (anonymous):

what % of 200 is 300

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

its more 200

OpenStudy (skullpatrol):

$$\Huge \frac {x}{100}=\frac{a}{b}$$

OpenStudy (triciaal):

percent means per 100 set up a ratio let the percentage = x 200/300 = x/100 solve for x

TheSmartOne (thesmartone):